# Client-wise Modality Selection for Balanced Multi-modal Federated Learning

### v1

@@ Abstract
Selecting proper clients to participate in the iterative federated learning (FL) rounds is critical to effectively harness a broad range of distributed datasets. Existing client selection methods simply consider the variability among FL clients with uni-modal data, however, have yet to consider clients with multi-modalities. We reveal that traditional client selection scheme in MFL may suffer from a severe modality-level bias, which impedes the collaborative exploitation of multi-modal data, leading to insufficient local data exploration and global aggregation. To tackle this challenge, we propose a Client-wise Modality Selection scheme for MFL (CMSFed) that can comprehensively utilize information from each modality via avoiding such client selection bias caused by modality imbalance. Specifically, in each MFL round, the local data from different modalities are selectively employed to participate in local training and aggregation to mitigate potential modality imbalance of the global model. To approximate the fully aggregated model update in a balanced way, we introduce a novel local training loss function to enhance the weak modality and align the divergent feature spaces caused by inconsistent modality adoption strategies for different clients simultaneously. Then, a modality-level gradient decoupling method is designed to derive respective submodular functions to maintain the gradient diversity during the selection progress and balance MFL according to local modality imbalance in each iteration. Our extensive experiments showcase the superiority of CMSFed over baselines and its effectiveness in multi-modal data exploitation.
@@ 1 Introduction
Figure 1: Traditional client selection in FL aims to sample a client subset in each round. The selected clients use all their local data for training and aggregation. To exploit different modalities comprehensively, we propose the client-wise modality selection for MFL, which considers the local data from each modality as the sampling unit rather than the client. Only the selected modality data from a client would join in training and aggregation.
@@
Federated learning (FL) [20] aims to collaboratively learn data that has been collected by, and resides on, a number of remote devices or servers. FL stands to develop top-performing models by aggregating knowledge from numerous edge clients [6], which relies on the iterative interaction among participating clients and the server. However, comprehensively employing the information from all clients can be exceptionally difficult due to the client heterogeneity and resource limitations [8, 27].
Random sampling [32, 19] from available clients is widely used in FL to satisfy some practical restrictions, e.g., limited communication bandwidth [23, 31] and computing capacities [13]. To improve the information exploitation of all clients, extensive research has been conducted on effective client selection strategies. One viable solution [24] is to gather as much local data as possible in each round while in contrast, some alternative studies [1] focus on collecting representative information by sampling a subset of clients. However, previous methods have only considered the knowledge discrepancies across different clients, while demonstrating notable performance in uni-modal FL, yet, their effectiveness diminishes when dealing with clients with multi-modal data as the inter-modality interactions during the FL training is neglected. The investigation into fine-grained and sophisticated data selection mechanism for multi-modal FL (MFL) is currently under-explored.
In this paper, we reveal that traditional client selection approaches may suffer from a severe modality-level bias in MFL. Due to the inconsistent learning progress of different modalities, i.e., the modality imbalance, client-level selection bias tends to favor information from a dominant modality while deprioritize the weak modality, which impedes the collaborative exploitation of distributed multi-modal data, leading to insufficient local data exploration and global aggregation.
To address such issue in MFL, we propose a novel Client-wise Modality Selection scheme for MFL (CMSFed), which chooses the appropriate modalities combination from clients in each round to alleviate the modality-level bias and realize full exploitation of each modality via considering the interactions between different modalities as illustrated in Fig. 1. Only part of, not all, modalities, will be selected for some clients compared with traditional client selection. Specifically, in each communication round, we intend to gather a data subset by selectively employing different modalities from each client, the aggregated model update stemmed from that approximates the fully aggregated update which fulfills the potential of each modality. Inspired from [1] for diverse client selection, we build the modality selection formula, namely the min-form of the facility location function [5] evaluated on the collected data subset, to seek for the representative information for the global model. To further release the ability of each modality and reduce the training cost, we propose a modality-level gradient decoupling with a modal-balanced local training loss to build two separated submodular functions and then modulate their optimization process for unbiased modality selection. As a result, inconsistent modality adoption strategies may lead to different feature spaces for different clients, which exacerbates the difficulties in accomplishing an accurate and generalized aggregated model. Therefore, the global prototype is introduced for the weak modality to alleviate the heterogeneous feature spaces. To summarize, our contributions in this paper are as follows:
  • We scrutinize the client selection scheme in MFL and reveal that naive sampling clients may introduce a strong modality-level bias which causes insufficient utilization of multi-modal data.
  • We propose a novel client-wise modality selection scheme for MFL (CMSFed) to comprehensively exploit all modalities via a enhancing loss and gradient decoupling to balance MFL.
  • We conduct comprehensive experiments and demonstrate that 1) CMSFed can achieve considerable improvements over baselines; 2) CMSFed indeed achieves a more balanced exploitation of distributed multimodal data.
@@ 2 Background and Related Works
(a) Multi-modal
(b) Uni-audio
(c) Uni-visual
Figure 2: Performance of client selection strategies on CREMA-D with uniformly distributed data on 20 clients. FedAvg randomly selects clients while DivFL selects clients with diversity. ‘FedAvg-0.8’ means 80% of clients have data in both audio and visual modalities, while 20% of clients only have data in random one modality. ‘-0.5’ and ‘-0.2’ follow the similar definition. (a) Performance of global multi-modal model. (b) Performance of global uni-audio model. (c) Performance of global uni-visual model.
@@ 2.1 Multi-modal federated learning
In MFL, each client may contain one or more modalities of data. Without loss of generality, we consider two input modalities, referred to as $A$ and $I$ in MFL. There are a set $V$ of $N$ clients, $V=\left[N\right]$, respectively owning datasets $\mathcal{D}_{i}=\left\{\boldsymbol{X}_{i}^{A},\boldsymbol{X}_{i}^{I},y\right\}$ composed of $n_{i}$ pair samples, $i\in\left[N\right]$. A typical federated learning objective is the average of each client’s local loss function:
$\vspace{-8pt}\min f\left(\theta\right)=\sum_{k=1}^{N}{p_{k}F_{k}\left(\theta\right)}$ (1)
where $\theta=\left\{\theta^{A},\theta^{I},\omega\right\}$ denotes the model parameters. $\theta^{A}$ and $\theta^{I}$ represent the parameters of encoders of modality $A$ and $I$. $\omega$ is the parameters of fusion classifier. $p_{k}$ is a pre-defined weight. $F_{k}$ is each client’s local loss.
@@ 2.2 Client selection and submodular function
$\sum_{k\in\left[N\right]}{\nabla F_{k}\left(v^{k}\right)}$ $=\sum_{k\in\left[N\right]}{\left[\nabla F_{k}\left(v^{k}\right)-\nabla F_{\sigma\left(k\right)}\left(v^{\sigma\left(k\right)}\right)\right]}$ (2)
$+\sum_{k\in S}{\gamma_{k}\nabla F_{k}\left(v^{k}\right)}$
@@
$\begin{split}&\left\|\sum_{k\in\left[N\right]}{\nabla F_{k}\left(v^{k}\right)}-\sum_{k\in S}{\gamma_{k}\nabla F_{k}\left(v^{k}\right)}\right\|\leqslant\\ &\sum_{k\in\left[N\right]}{\underset{i\in S}{\min}\left\|\nabla F_{k}\left(v^{k}\right)-\nabla F_{i}\left(v^{i}\right)\right\|\triangleq G\left(S\right)}\end{split}$ (3)
@@
$S\leftarrow S\cup k^{*},$ (4)
$k^{*}\in$ $\underset{k\in\mathrm{rand}\left(V\backslash S,\mathrm{size}=s\right)}{\mathrm{arg}\max}\left[\bar{G}\left(S\right)-\bar{G}\left(\left\{k\right\}\cup S\right)\right]$
@@
Although these methods make great improvement in uni-modal FL, the selection strategy is under-explored in MFL and we reveal that traditional client selection approaches fall into a severe bias in MFL.
@@ 2.3 Imbalanced multi-modal learning
3 Observations and Analysis
In this section, we analyze the behavior of different client sampling strategies in MFL and the interactions between different modalities. As shown in Fig. 2, we compare a well-designed client selection method (DivFL) with random sampling (FedAvg) on a multi-modal dataset. The results in Fig. 1(a) show that the improvement from DivFL is severely limited although it is terrific in uni-modal FL. The results about how each modality behaves are illustrated in Figs. 1(b) and 1(c). It is apparent that audio modality performs much better than visual modality and the optimised client selection options from DivFL generally focus on audio modality (audio performance of DivFL has significant improvement over FedAvg while no noticeable difference in visual).
The performance gap between audio and visual can be attributed to ‘modality imbalance’ [33], which indicates the inconsistent learning paces in different modalities. Therefore, the imbalance nature (multi-modal learning is dominated by the modality who learns fast) [34] in MFL leads to a strong modality-level selection bias, i.e., client selection strategies tend to select clients that make the dominant modality aggregation better but neglect the weak modality
To further analyze the interactions between different modalities, we randomly discard the data of one modality (audio or visual) on clients with certain probabilities (FedAvg-0.8, -0.5, -0.2) as demonstrated in Fig. 1(a). It is clear that randomly discarding one modality data for part of clients does not necessarily lead to performance degradation of the global model, but may bring performance improvement. The reason behind can be found in Figs. 1(b) and 1(c), where the gap between audio and visual modalities gradually narrows as the discarding probability increases until there is no stark disparity (in FedAvg-0.2). This phenomenon is consistent with the findings in [26]: the dominant modality hinders the learning process of weak modality. Therefore, the ability of visual modality is released and absorbed by the global model through aggregation when the audio data in some clients is abandoned.
The above analyses suggest that traditional client selection methods are not applicable in MFL due to the modality imbalance nature, and the potential of each modality can be further explored by engaging appropriate modalities in training. Therefore, we intend to design a new modality selection scheme in MFL to address the modality imbalance and information underutilization from different modalities.
@@ 4 Method
In this section, we introduce CMSFed, a diverse and balanced modality selection strategy to realize more efficient data utilization. We first derive a combinatorial objective for modality selection in MFL via the gradient approximation with the fully aggregated update. Then we propose a modality-level gradient decoupling with the enhancing gradient on weak modality to build two separated submodular functions and balance MFL.
4.1 Client-wise modality selection
Firstly, we rewrite Eq. 2 for MFL as:
$\begin{split}&\sum_{k\in\left[N\right]}{\nabla F_{k}\left(v_{A}^{k},v_{I}^{k}\right)}=\sum_{k\in S}{\gamma_{k}\nabla F_{k}\left(v_{A}^{k},v_{I}^{k}\right)}+\\ &\sum_{k\in\left[N\right]}{\left[\nabla F_{k}\left(v_{A}^{k},v_{I}^{k}\right)-\nabla F_{\sigma\left(k\right)}\left(v_{A}^{\sigma\left(k\right)},v_{I}^{\sigma\left(k\right)}\right)\right]}\end{split}$ (5)
where $v_{A}^{k}$ and $v_{I}^{k}$ indicate the corresponding modal data in client $k$.
The above equation is still the client selection formulation, while we could build the modality selection formula based on it as (data superscripts are omitted for simplicity):
$\sum_{k\in\left[N\right]}{\nabla F_{k}\left(v_{A},v_{I}\right)}$ (6)
$=\sum_{k\in\left[N\right]}{\left[\begin{array}[]{c}\nabla F_{k}\left(v_{A},v_{I}\right)-\nabla F_{\sigma_{M}\left(k\right)}\left(v_{A},v_{I}\right)\\ -\nabla F_{\sigma_{A}\left(k\right)}\left(v_{A}\right)-\nabla F_{\sigma_{I}\left(k\right)}\left(v_{I}\right)\\ \end{array}\right]}$
$+\sum_{k\in S_{M}}{\gamma_{k}^{M}\nabla F_{k}\left(v_{A},v_{I}\right)}+\sum_{k\in S_{A}}{\gamma_{k}^{A}\nabla F_{k}\left(v_{A}\right)}$
$+\sum_{k\in S_{I}}{\gamma_{k}^{I}\nabla F_{k}\left(v_{I}\right)}$
where $\sigma_{M}$, $\sigma_{A}$ and $\sigma_{I}$ map $V\rightarrow S_{M},S_{A},S_{I}$, the client sets who use multi-modal, $A$ modal, and $I$ modal data for training respectively, and $S_{M}\cap S_{A}=S_{A}\cap S_{I}=S_{M}\cap S_{I}=\oslash$. The full gradient aggregation is represented by the customized update from multi-modal and uni-modal clients. Assume modality $I$ is weak here, the relaxed objective for modality selection should be:
$\left\|\begin{array}[]{c}\sum_{k\in\left[N\right]}{\nabla F_{k}\left(v_{A},v_{I}\right)}-\sum_{k\in S_{I}}{\gamma_{k}^{I}\nabla F_{k}\left(v_{I}\right)}\\ -\sum_{k\in S_{M}}{\gamma_{k}^{M}\nabla F_{k}\left(v_{A},v_{I}\right)}\\ \end{array}\right\|$ (7)
$\leqslant\sum_{k\in\left[N\right]}{\underset{i\in S_{M},j\in S_{I}}{\min}\left\|\begin{array}[]{c}\nabla F_{k}\left(v_{A},v_{I}\right)-\\ \nabla F_{i}\left(v_{A},v_{I}\right)-\nabla F_{j}\left(v_{I}\right)\\ \end{array}\right\|}$
uni-$A$ modal is omitted here as the multi-modal gradient is dominated by $A$, which means we do not need to select uni-$A$ clients for its acceleration. However, minimizing this objective, which is similar to client selection in Eq. 3, cannot leverage sufficient information from each modality since the approximation objective (the left-hand side of Eq. 6) still suffers from serious modality imbalance. How to finding a new full gradient aggregation that can exploit all modalities’ information is a great challenge. In addition to this, there also exist other challenges making Eq. 7 difficult to be effectively optimized in practice:
  • We need to build the relationship (a gradient similarity matrix) between fully aggregated gradient and local gradient for solving Eq. 7. In uni-modal FL, a client only needs to train once to get gradients while in MFL, each client should perform local update twice (with multi-modal and weak-modal data respectively) to determine whether to choose multi-modal or uni-modal data for training. This inevitably leads to almost double the local training cost.
  • Applying modality selection would cause different modality adoption strategies, namely modality heterogeneity, which results in divergent feature spaces across clients, especially on weak modality.
To address the above challenges, we propose a modality-level gradient decoupling method with a balanced local loss, which not only leads to comprehensive information exploitation but also notably saves training overhead.
4.2 Modality-level gradient decoupling
We focus on the classification tasks in this paper, where cross-entropy (CE) loss is widely used. Hence, the loss for multi-modal and uni-modal learning should be:
$\mathrm{multi}\mbox{-}\mathrm{modal}$ $:F_{k}\left(v_{A},v_{I}\right)=\mathcal{L}_{CE}^{k}\left(v_{A},v_{I}\right)$ (8)
$\mathrm{uni}\mbox{-}\mathrm{modal}$ $:F_{k}\left(v_{I}\right)=\mathcal{L}_{CE}^{k}\left(v_{I}\right)$
The naïve multi-modal loss falls into modality imbalance. Although obtaining a totally balanced solution seems impractical for multi-modal learning, we can find a surrogate for it: here we enhance the weak modality by equipping the naïve losses with the prototypical cross-entropy (PCE) loss $\beta^{k}\mathcal{L}_{P}^{k}$ [9] to correct gradient directions of weak modality:
$\mathrm{multi}\mbox{-}\mathrm{modal}$ $:\tilde{F}_{k}\left(v_{A},v_{I}\right)=\mathcal{L}_{CE}^{k}\left(v_{A},v_{I}\right)+\beta^{k}\mathcal{L}_{P}^{k}\left(v_{I}\right)$ (9)
$\mathrm{uni}\mbox{-}\mathrm{modal}$ $:\tilde{F}_{k}\left(v_{I}\right)=\mathcal{L}_{CE}^{k}\left(v_{I}\right)+\beta^{k}\mathcal{L}_{P}^{k}\left(v_{I}\right)$
where $\mathcal{L}_{P}^{k}$ is calculated based on the local prototypes $c^{I}$:
$\mathcal{L}_{P}^{k}\left(v_{I}\right)=-\mathbb{E}_{\left(x_{i}^{I},y\right)\in\mathcal{D}_{k}}\log\left[\frac{\exp\left(-d\left(z_{i}^{I},c_{y}^{I}\right)\right)}{\sum\nolimits_{j=1}^{Y}{\exp\left(-d\left(z_{i}^{I},c_{j}^{I}\right)\right)}}\right]$ (10)
where $d\left(\cdot,\cdot\right)$ is the distance function (Euclidean distance), $z_{i}^{I}$ is the feature of $x_{i}^{I}$. $Y$ is the class number. The calculation details for $c^{I}$, $\beta^{k}$ are in Appendix. Applying $\mathcal{L}_{P}^{k}$ on uni-modal loss is also essential, which can improve this modality by clustering features tightly.
Bring Eq. 9 to Eq. 7 and we can get the new objective:
$\sum_{k\in\left[N\right]}{\underset{i\in S_{M},j\in S_{I}}{\min}\left\|\begin{array}[]{c}\nabla\tilde{F}_{k}\left(v_{A},v_{I}\right)-\\ \nabla\tilde{F}_{i}\left(v_{A},v_{I}\right)-\nabla\tilde{F}_{j}\left(v_{I}\right)\\ \end{array}\right\|}=$ (11)

### v2

# Overcome Modal Bias in Multi-modal Federated Learning via Balanced Modality Selection

@@ Abstract
Selecting proper clients to participate in each federated learning (FL) round is critical to effectively harness a broad range of distributed data. Existing client selection methods simply consider the mining of distributed uni-modal data, yet, their effectiveness may diminish in multi-modal FL (MFL) as the modality imbalance problem not only impedes the collaborative local training but also leads to a severe global modality-level bias. We empirically reveal that local training with a certain single modality may contribute more to the global model than training with all local modalities. To effectively exploit the distributed multiple modalities, we propose a novel Balanced Modality Selection framework for MFL (BMSFed) to overcome the modal bias. On the one hand, we introduce a modal enhancement loss during local training to alleviate local imbalance based on the aggregated global prototypes. On the other hand, we propose the modality selection aiming to select subsets of local modalities with great diversity and achieving global modal balance simultaneously. Our extensive experiments on audio-visual, colored-gray, and front-back datasets showcase the superiority of BMSFed over baselines and its effectiveness in multi-modal data exploitation.
Keywords:
Multi-modal federated learning modality imbalance modality selection
@@ 1 Introduction
Federated learning (FL) [23] aims to collaboratively learn data that has been collected by, and resides on, a number of remote devices or servers. FL stands to develop top-performing models by aggregating knowledge from numerous edge clients [8, 36], which relies on the iterative interaction among participating clients and the server. However, comprehensively employing the information from all clients can be exceptionally difficult due to the client heterogeneity and resource limitations [10, 29].
Random sampling [34, 22] from available clients has been widely used in FL to satisfy some practical restrictions, e.g., limited communication bandwidth [26, 35] and computing capacities [15]. To improve the information exploitation of all clients, extensive research has been conducted on effective client selection strategies [9, 43]. Despite the success of traditional client selection methods in uni-modal FL, their effectiveness diminishes when dealing with clients with multi-modal data as the inter-modal interactions during the MFL training are neglected. According to [37, 14], there may exist inconsistent learning paces for different modalities in multi-modal joint training, i.e., modality imbalance, which not only impedes the collaborative local training but also leads to a severe modal bias for global model in MFL. As illustrated in Tab. 1, audio modality significantly outperforms visual modality in CREMA-D and AVE datasets during local training and the aggregated model still suffers from it. However, two well-designed client selection methods (pow-d [6] and DivFL [3]) only obtain severely limited improvement over random sampling (FedAvg) on the multi-modal global model. We can see that client selection methods achieve the best uni-audio performance while uni-visual performance even drops sometimes, which means existing client selection scheme heavily relies on the better modality, while ignoring the importance of improving weak modalities that also has potential for global model aggregation. Based on the above analysis, a pivotal question arises: Can we design a new selection scheme in MFL that can overcome the modal bias and exploit each modality comprehensively?
Table 1: Performance of various client selection methods in MFL under IID setting. A and V denote uni-audio and uni-visual while A-V means the multi-modal result. ‘Local’ represents that a client is trained based its local data without aggregation. A strong modal bias of global model exists on the two datasets.
Dataset CREMA-D [4] AVE [33]
Method A V A-V A V A-V
Local 41.9 20.4 39.6 33.4 16.7 35.2
FedAvg 51.2 20.6 50.7 61.1 26.8 62.2
[…]
DivFL [3] 52.3 21.1 51.7 62.7 25.3 63.3
FedAvg-0.2 50.6 28.6 52.4 60.6 29.6 63.4
FedAvg-0.5 50.5 34.6 55.7 58.7 30.0 60.7
FedAvg-0.8 48.1 50.9 61.2 56.4 31.8 58.5
BMSFed 51.0 41.9 64.5 59.7 40.2 64.7
To answer this question, we investigate the interactions between different modalities via randomly discarding the data from one modality (audio or visual) on part of clients, which is inspired from modality dropout [2, 40] that drops a specific modal data during training for regularization. The results are shown in rows 7-9 in Tab. 1, where ‘-x’ denotes randomly discarding a modality on a client with probability ‘x’. We can see that dropping with a certain probability can improve the global multi-modal performance on both datasets (e.g., FedAvg-0.2), and the main reason comes from the dramatic improvement of visual modality. This phenomenon suggests that performing uni-modal training can unleash its potential without being inhibited by another modality, and uni-modal local training may contribute more to the global model than multi-modal training on some clients . As the dropping ratio increases, although the visual modality still improves, multi-modal performance may decline as the audio modality declines as shown in columns 5 and 7, suggesting that we should carefully control which modalities on each client should be involved in training and aggregation to make the contribution most to the global model.
Figure 1: Left: Traditional client selection in FL aims to sample a client subset in each round while our modality selection considers each local modality as the sampling unit. Right: The paradigm of BMSFed with four clients. The global prototypes are used to enhance the weak modality during local update. Only networks corresponding to the selected modalities will be uploaded to the server for aggregation.
According to the above investigations, we propose a novel Balanced Modality Selection scheme for MFL (BMSFed) to mitigating the modal bias and comprehensively exploit the diverse information from all modalities. Specifically, instead of selecting a subset of clients, we treat each modality on the local side as a selection unit, as demonstrated in Fig. 1. Our BMSFed mainly contains two parts: Firstly, we intend to alleviate the local modality imbalance by introducing a modal enhancement loss based on aggregated global prototypes to promote the performance of weak modality. Secondly, we complete the modality selection by building two separated submodular functions for selecting multi-modal clients (training with multi-modal data) and uni-modal clients (training with selected uni-modal data) respectively. Inspired from [3], the criterion is to select modalities that are most representative on the gradients while also alleviate the global modal bias. A simple yet effective conflict resolution strategy is devised to ensure the validity of modality selection and keep modal balance on global model simultaneously.
The main contributions of the paper are summarized as follows:
  1. 1.
    We empirically analyze the modality imbalance problem in MFL and reveal that uni-modal training on some clients may contribute more to the global model than multi-modal training.
  2. 2.
    Based on the analysis, we propose a novel Balanced Modality Selection scheme for MFL (BMSFed) to comprehensively exploit all modalities via a modal enhancement loss and representative modality selection to overcome the global modal bias.
  3. 3.
    We conduct comprehensive experiments on audio-visual, colored-gray, and front-back datasets, and considering the statistical heterogeneity and modality incongruity problems in MFL, to validate the superiority of our BMSFed.
@@ 2.1 Multi-modal federated learning
In MFL, each client has one or multiple modalities of data. Without loss of generality, we consider two input modalities, which are denoted by $A$ and $I$ respectively in MFL. There are a set $V$ of $N$ clients, $V=\left[N\right]$, respectively owning datasets $\mathcal{D}_{i}=\left\{\boldsymbol{X}_{i}^{A},\boldsymbol{X}_{i}^{I},\boldsymbol{y}\right\}$, $i\in\left[N\right]$. A typical federated learning objective is the average of each client’s local loss function:
$\vspace{-10pt}\min f\left(\theta\right)=\sum_{k=1}^{N}{p_{k}F_{k}\left(\theta\right)}$ (1)
where $\theta=\left\{\theta^{A},\theta^{I},\omega\right\}$ denotes the model parameters. $\theta^{A}$ and $\theta^{I}$ represent the encoder parameters of modality $A$ and $I$. $\omega$ is the parameter of fusion classifier. $p_{k}$ is a pre-defined weight. $F_{k}$ is each client’s local loss (cross entropy (CE) loss for classification task in this paper).
@@ 2.2 Client selection and submodular function
$\sum_{k\in\left[N\right]}{\nabla F_{k}\left(v^{k}\right)}=\sum_{k\in\left[N\right]}{\left[\nabla F_{k}\left(v^{k}\right)-\nabla F_{\sigma\left(k\right)}\left(v^{\sigma\left(k\right)}\right)\right]}+\sum_{k\in S}{\gamma_{k}\nabla F_{k}\left(v^{k}\right)}$ (2)
@@
$\begin{split}\left\|\sum_{k\in\left[N\right]}{\nabla F_{k}\left(v^{k}\right)}-\sum_{k\in S}{\gamma_{k}\nabla F_{k}\left(v^{k}\right)}\right\|\leqslant\sum_{k\in\left[N\right]}{\underset{i\in S}{\min}\left\|\nabla F_{k}\left(v^{k}\right)-\nabla F_{i}\left(v^{i}\right)\right\|\triangleq G\left(S\right)}\end{split}$ (3)
@@
$S\leftarrow S\cup k^{*},k^{*}\in\underset{k\in\mathrm{rand}\left(V\backslash S,\mathrm{size}=s\right)}{\mathrm{arg}\max}\left[\bar{G}\left(S\right)-\bar{G}\left(\left\{k\right\}\cup S\right)\right]$ (4)
@@
Although these methods make great improvement in uni-modal FL, the selection strategy is under-explored in MFL and we reveal that traditional client selection approaches cannot address the severe modal bias in MFL.
@@ 3 Method
In this section, we introduce BMSFed that contains the local imbalance alleviation and balanced modality selection.
3.1 Local Imbalance Alleviation
As discussed in Sec. 2.3 and Tab. 1, the multi-modal training on each client may suffer from severe modality imbalance, leading to inadequate information exploitation and consequently incurring the modal bias on the global model. Therefore, we first try to alleviate the imbalance during local training.
According to [11], the prototypes (i.e., class centroids) are suitable to calibrate the gradient directions to avoid the interference from other modalities. Therefore, we choose to use prototypes to facilitate learning of weak modality. For the $i$-th client with the data $\mathcal{D}_{i}=\left\{\boldsymbol{X}_{i}^{A},\boldsymbol{X}_{i}^{I},\boldsymbol{y}\right\}$, the local prototype for class $j$ of each modality is defined as the mean value of representations:
$c_{i,j}^{I}=\frac{1}{\left|\mathcal{D}_{i,j}\right|}\sum_{x^{I}\in\mathcal{D}_{i,j}}{h_{i}\left(\theta_{i}^{I};x^{I}\right)},c_{i,j}^{A}=\frac{1}{\left|\mathcal{D}_{i,j}\right|}\sum_{x^{A}\in\mathcal{D}_{i,j}}{h_{i}\left(\theta_{i}^{A};x^{A}\right)}$ (5)
where $\mathcal{D}_{i,j}$ denotes the samples belonging to $j$-th class in client $i$. $h_{i}$ is the function of encoder. Considering the heterogeneity across clients, we aggregate the local prototypes to a global prototype as:
$c_{j}^{GI}=\frac{1}{\left|\mathcal{N}_{j}\right|}\sum_{i\in\mathcal{N}_{j}}{\frac{\left|\mathcal{D}_{i,j}\right|}{N_{j}}c_{i,j}^{I}},c_{j}^{GA}=\frac{1}{\left|\mathcal{N}_{j}\right|}\sum_{i\in\mathcal{N}_{j}}{\frac{\left|\mathcal{D}_{i,j}\right|}{N_{j}}c_{i,j}^{A}}$ (6)
where $\mathcal{N}_{j}$ denotes the set of clients that have class $j$ and $N_{j}$ is the number of instances belonging to class $j$ over all clients. Then, we introduce a modal enhancement loss (ME) based on global prototype to adjust local training:
$\mathcal{L}_{ME}^{k}\left(v_{I}^{k}\right)$ $=-\mathbb{E}_{\left(x_{i}^{I},y\right)\in\mathcal{D}_{k}}\log\left[\frac{\exp\left(-d\left(z_{i}^{I},c_{y}^{GI}\right)\right)}{\sum\nolimits_{j=1}^{Y}{\exp\left(-d\left(z_{i}^{I},c_{j}^{GI}\right)\right)}}\right]$ (7)
$\mathcal{L}_{ME}^{k}\left(v_{A}^{k}\right)$ $=-\mathbb{E}_{\left(x_{i}^{A},y\right)\in\mathcal{D}_{k}}\log\left[\frac{\exp\left(-d\left(z_{i}^{A},c_{y}^{GA}\right)\right)}{\sum\nolimits_{j=1}^{Y}{\exp\left(-d\left(z_{i}^{A},c_{j}^{GA}\right)\right)}}\right]$
where $d\left(\cdot,\cdot\right)$ is the distance function (Euclidean distance), $z_{i}^{I}$ is the representation of $x_{i}^{I}$, i.e., $z_{i}^{I}=h_{i}\left(\theta_{i}^{I};x_{i}^{I}\right)$. $Y$ is the class number. $v_{A}^{k}$ and $v_{I}^{k}$ indicate the corresponding modal data in client $k$. Hence, the local loss should be (data superscripts are omitted for simplicity):